OpenStudy (anonymous):

-3x+6y=12 in standard form

OpenStudy (anonymous):

i dont exactly remember how it goes but you make it so it ends up like this x = y or y = x

OpenStudy (anonymous):

Standard form is this : y=mx+d In there we should change the place of -3x (mx) . Can you do that ?:)

OpenStudy (anonymous):

yes thanks alot

OpenStudy (anonymous):

Got it ?:)????????!

OpenStudy (anonymous):

But it isnt enough !:)

OpenStudy (anonymous):

Then we have : 6y=3x+12 and should be kill 6 for y . So have : y=3/6x+2 Right ? And can say ; y=1/2x+2 Got it ?:)